Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Brief answer to selected problems in HW01

1. Section 2.3: Problem 35. Take δ = min (3 + \frac{3.5^{2}−1}{−5}, −(3 + \frac{4.5^{2}−1}{−5})). It will work for ε = 0.5.

2. Section 2.3: Problem 49. Taking δ = ε will do.

3. Section 2.3: Problem 53. The function f(x) = x^{2}, if x ≠ 0 with f(0) = −1 satisfies the statement with x_{0} = 0 and L = −1 but the limit is not L.